40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 18 12:40:09 2019, gautam, Update, CDS, Fast BIO Mapping at 1Y2 
    Reply  Thu Sep 19 21:23:51 2019, gautam, Update, CDS, Fast BIO splicing re-implemented at 1Y2 1Y2_FAST_BIO_WIRING_MAP.pdfIMG_7949.JPGREFL165.pdf
Message ID: 14901     Entry time: Thu Sep 19 21:23:51 2019     In reply to: 14895
Author: gautam 
Type: Update 
Category: CDS 
Subject: Fast BIO splicing re-implemented at 1Y2 

[KA, GV]

Summary:

  1. New cross connect system for splicing the fast BIO signals for whitening switching to the P2 connectors was installed and tested at 1Y2.
  2. It passed a first round of tests. 😁 
  3. As of now, I believe all the necessary electrical connections have been made at 1Y2/1Y3, and we are ready for testing the c1iscaux system.

Details:

  1. We did some testing in the office area, and found several wiring mistakes. These were all rectified. Attachment #1 is an accurate reflection of the implemented wiring scheme (softcopy in the 40m google sheets area). Be aware that the IDC 50 pin connector pin-out is tricky, and you have to be aware of the difference between male/female connector when looking for this pin-out on the internet.
  2. In order to facilitate further testing, we re-routed the ADC0 SCSI cable that was unplugged on the overhead cable tray, and plugged it back into the c1lsc expansion chassis. This action necessitated a reboot of the vertex FEs, but everything came back alright.
  3. Did some general neatenign and strain relieving. Removed a few existing cross-connects to make space for our new terminal blocks.
  4. Attachment #2 shows the layout of the terminal blocks. Note the unusual (vertical) order of the orange terminal blocks.
  5. The final integrated CDS test done was the following:
    • Set whitening gain for channel under test to 45dB, so that the dark noise level is boosted to a measurable level such that a change can be seen with the whitening enabled/disabled.
    • Compare the ASD of the signal between 30-100 Hz with the whitening engaged/disengaged.
    • Example result shown in Attachment #3.I believe the whitening is 15:150 (z:p) 

Tomorrow:

  1. Recover POX/POY locking,.
  2. ...
Quote:

Update 2019 Sep 19 1730: The pin numbers of the IDC 50 connector are all off by 1. i.e. 3-->4 and so on. I will fix this shortly. The problem was because of me looking at the pinout for the wrong gender of IDC50 connectors.

Attachment 1: 1Y2_FAST_BIO_WIRING_MAP.pdf  46 kB  | Hide | Hide all
1Y2_FAST_BIO_WIRING_MAP.pdf
Attachment 2: IMG_7949.JPG  2.634 MB  Uploaded Thu Sep 19 22:34:26 2019  | Hide | Hide all
IMG_7949.JPG
Attachment 3: REFL165.pdf  33 kB  Uploaded Thu Sep 19 23:54:49 2019  | Hide | Hide all
REFL165.pdf
ELOG V3.1.3-